(1 point) Find the constant term in the expansion of (-1x7 + 48 The constant term is(1 point) Find the constant term in the expansion of (?1x7+?3x)48(?1x7+?3x)48.
Answer #1

General term of the expansion,

Tr = 48Cr (-x^7)^(48-r) (-3/x)^r

= 48Cr x^7(48-r)(-3)^r(x)^-r

= 48Cr x^(336-7r-r) (-3)^r

= 48Cr x^(336-8r) (-3)^r

For constant term, power of x should be zero,so,

336 - 8r = 0 => r = 336/8 = 42

Hence 43 rd term will be a constant term.